b1az34:

Write the slope-intercept form of the equation of the line that is perpendicular to AB and passes through Point X. Show all work for full credit

b1az34:

1 attachment
kacchan:

' y=(-5/12)x+8 ' find y=mx+b for line parallel to AB. the slope of line AB is m=(8-3)/(-10-2) =5/-12 =-5/12

kacchan:

i believe you would get y=(-5/12)x+8 as your answer

Eveeeeeeeee14:

wow you are good at math angle

kacchan:

@eveeeeeeeee14 wrote:
wow you are good at math angle
ik

Eveeeeeeeee14:

I'm not good at math at all lol

kacchan:

@eveeeeeeeee14 wrote:
I'm not good at math at all lol
oof

snowflake0531:

respectfully disagree with previous answer first, the equation of AB the slope is -5/12 to find the y -intercept, take the slope and one coordinate (2,3) 3 = -5/12 (2) + b 3 = -5/6 + b b = 3 5/6 so the equation of AB is \(y=−\frac{5}{12}x+3\frac{5}{6}\) Now, we have to find the line perpendicular, that goes through (-5,10) let's first find the slope the perpendicular line's slope is always negative reciprocal, so the slope would be 12/5 now, we use the slope and coordinate to find the y intercept 10 = 12/5 (-5) + b 10 = -12 + b b = 22 so the perpendicular line is \(y=\frac{12}{5}x+22\)

b1az34:

@snowflake0531 u saved me thx

snowflake0531:

yw~